OpenStudy (anonymous):

Greg has 4 shirts: a white one, a black one, a red one, and a blue one. He also has two pairs of pants, one blue and one tan. What is the probability, if Greg gets dressed in the dark, that he winds up wearing the white shirt and tan pants? Show your work.

OpenStudy (anonymous):

1/6th

OpenStudy (anonymous):

will medal and fan

OpenStudy (kropot72):

The probability of choosing a white shirt is 1/4 and the probability of choosing the tan pants is 1/2. The two events are independent, therefore the probability of 'white shirt/tan pants' is given by: \[\large P(white\ shirt,\ \tan\ pants)=\frac{1}{4}\times\frac{1}{2}=you\ can\ calculate\]